A patient has an attendant that delivers 2.70 tsp of Elixir Sulfanilamide 5 times per day. If the elixir is essentially 100% DEG, what is the total mass of DEG delivered to the patient in a 24 hour period?
Answer #1

1 tsp = 5 grams

2.7 tsp = 5*2.7 = 13.5 g

She serves 5 times per day

no. of grams per day = 13.5 * 5 = 67.5 g of DEG

So, in a 24 hour period she serves 2.7*5 = 13.5 tsp or 67.5 g of DEG.